A ball is thrown vertically upward with a speed of 7.1 m/s. Find:




(9 points)




a. Maximum height reached the ball




b. Its time of flight




c. Velocity the ball would return to its starting point

1
Expert's answer
2022-05-09T09:45:05-0400
"v(t)=v_0-gt"

"h(t)=h_0+v_0t-\dfrac{gt^2}{2}"

a. The ball reaches a maximum height at


"t_1=\dfrac{v_0}{g}"

"h_{max}=h_0+\dfrac{v_0^2}{2g}"

Given "h_0=0m, v_0=7.1m\/s, g=9.81m\/s^2"


"h_{max}=0m+\dfrac{(7.1m/s)^2}{2(9.81m/s^2)}=2.569m"

b.


"t_{flight}=2t_1=\dfrac{2v_0}{g}"

"t_{flight}=\dfrac{2(7.1m/s)}{9.81m/s^2}=1.4475s"

c.


"v_{return}=v_0-gt_{flight}=-v_0"

"v_{return}=-7.1m\/s"